OpenStudy (anonymous):

Determine the domain, range, and horizontal asymptote of f(x)=(3/4)^(x+3)

OpenStudy (campbell_st):

the asymptote is y = 0.... as x gets extremely large (3/4)^(x+3) gets close to zero.... so the domain is all real x.... same as any exponential function range is y> 0

Directrix (directrix):

Domain: set of Reals Range: y > 0 Horizontal Asymptote: y = 0

OpenStudy (anonymous):

How is it I go in and find the real #s....would it be 3/4

OpenStudy (anonymous):

I need to know how to work all of it out

OpenStudy (anonymous):

since there is no denominator, domain is always real # if there is a denominator then be careful those values which makes denominator to zero..

OpenStudy (anonymous):

I dont know how to put any of it into an equation or whatever to know what is what, meaning real #s

OpenStudy (anonymous):

for example x^2/(x-2) domain all real # except 2

OpenStudy (anonymous):

so in this equation my real #s would be 3/4 and 3

OpenStudy (anonymous):

N={0,1,2,...... }naturel #s Z={.......-1,0.1,2,....} integers Q=a/b a and b Z R=Q+ Irrational # i.e Q= 1/2, -3/23, 2 , 0 24/1000 I=sqrt3, Pi, 1/2+srtq8 R={.....-1000.....-2,-1,0,1/2,3/2,sqrt 28.......}

OpenStudy (anonymous):

\[R=(-\infty, \infty)\]

OpenStudy (anonymous):

real #s any numbers except complex numbers..

OpenStudy (anonymous):